### v1

Abstract
We continue to study the learning-theoretic foundations of generation by extending the results from Kleinberg and Mullainathan (2024) and Li et al. (2024) to account for noisy example streams. In the noiseless setting of Kleinberg and Mullainathan (2024) and Li et al. (2024), an adversary picks a hypothesis from a binary hypothesis class and provides a generator with a sequence of its positive examples. The goal of the generator is to eventually output new, unseen positive examples. In the noisy setting, an adversary still picks a hypothesis and a sequence of its positive examples. But, before presenting the stream to the generator, the adversary inserts a finite number of negative examples. Unaware of which examples are noisy, the goal of the generator is to still eventually output new, unseen positive examples. In this paper, we provide necessary and sufficient conditions for when a binary hypothesis class can be noisily generatable. We provide such conditions with respect to various constraints on the number of distinct examples that need to be seen before perfect generation of positive examples. Interestingly, for finite and countable classes we show that generatability is largely unaffected by the presence of a finite number of noisy examples.
1 Introduction
Generation is an important paradigm in machine learning with promising applications to natural language processing (Wolf et al., 2020), computer vision (Khan et al., 2022), and computational chemistry (Vanhaelen et al., 2020). In contrast to its practical applications, a strong theoretical foundation of generation is largely missing from literature. Recently, Kleinberg and Mullainathan (2024), inspired by the seminal work of E Mark Gold on language identification in the limit (Gold, 1967), introduced a theoretical model of language generation called “Language Generation in the Limit.” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subset U$ for all $i\in\mathbbm{N}$. An adversary picks a language $K\in C$, and begins to enumerate the strings one by the one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses a string $\hat{w}_{t}\in U$ with the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. The class $C$ is generatable in the limit, if the player can generate from all $K\in C.$ Unlike language identification in the limit, Kleinberg and Mullainathan (2024) prove that generation in the limit is possible for every countable language family $C$.
Kleinberg and Mullainathan (2024)’s positive result has spawned a surge of new work, extending the results of Kleinberg and Mullainathan (2024) in various ways (Kalavasis et al., 2024b; Li et al., 2024; Charikar and Pabbaraju, 2024; Kalavasis et al., 2024a). Of particular interest to us is the work by Li et al. (2024). They frame the results of Kleinberg and Mullainathan (2024) through a learning-theoretic lens by taking the language family $C$ to be a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over some countable instance space $\mathcal{X}.$ By doing so, Li et al. (2024) identify stronger notions of generatability, termed “uniform”  [^1] Uniform generatability was also informally considered by Kleinberg and Mullainathan (2024) and “non-uniform” generatability, and provide characterizations of which classes are uniformly and non-uniformly generatable in terms of new combinatorial dimensions.
A commonality of Kleinberg and Mullainathan (2024) and its follow-up work is the assumption that the adversary presents to the player a “noiseless” stream of examples – one where every example/string must be contained in the hypothesis/language chosen by the adversary. In practice, such an assumption is unrealistic, as one would still like to generate well even if a few examples in the dataset are imperfect. For example, in light of recent interest in adversarial robustness, one might want a generative model to be robust to data contamination/poisoning attacks (Shang et al., 2018; Zhang et al., 2023; Jiang et al., 2023).
Motivated by these concerns, we study generation, in the framework posed by Kleinberg and Mullainathan (2024) and Li et al. (2024), under noisy example streams. In particular, we focus on a simple, but natural noising process: after selecting a positive example stream, the adversary is allowed to insert a finite number of positive examples in any way it likes, unbeknownst to the player. Such a noising process, as well as others, have been extensively studied in the context “language identification in the limit” or “inductive inference” (Baliga et al., 1992; Fulk and Jain, 1989; Jain, 1994; Schäfer, 1985; Stephan, 1997; Mukouchi and Sato, 2003; Tantini et al., 2006; Lange and Grieser, 2002; Case et al., 1997). In this paper, we extend this study to generation. To that end, our main contributions are summarized below.
  • (1)
    We extend the notions of uniform generatability, non-uniform generatability, and generatability in the limit from Kleinberg and Mullainathan (2024) and Li et al. (2024) to account for finite, noisy example streams. We call these new settings “noisy uniform generatability,” “noisy non-uniform generatability,” and “noisy generatability in the limit” respectively.
  • (2)
    We provide a complete characterization of which classes are noisily uniformly generatable in terms of a new scale-sensitive dimension we call the Noisy Closure dimension.
    Theorem (Informal).
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily uniformly generatable if and only if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for every $n\in\mathbb{N}$, where $\operatorname{NC}_{n}(\mathcal{H})$ is the Noisy Closure dimension of $\mathcal{H}$ at noise-level $n.$
    We show that noisy uniform generation is strictly harder than noiseless uniform generation. In fact, we construct an countably infinite class which is trivially uniformly generatable when there is no noise, but not noisily uniformly generatable even when the adversary is only allowed to perturb a single example. Despite this hardness, we show that all finite classes are still noisily uniformly generatable.
  • (3)
    We provide a sufficient condition for which classes are noisily non-uniformly generatable.
    Lemma (Informal).
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily non-uniformly generatable if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    Using the result that all finite classes are noisily uniformly generatable, we prove as a corollary that all countable classes are noisily non-uniformly generatable. Since noisy non-uniform generation implies noisy generation in the limit, this corollary also implies that all countable classes are noisily generatable in the limit. Although not matching, we also provide a necessary condition for noisy non-uniform generation in terms of the Noisy Closure dimension.
    Lemma (Informal).
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily non-uniformly generatable only if for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    We leave the complete characterization of noisy non-uniform generatability as an open question.
  • (4)
    We provide two different sufficiency conditions for noisy generatability in the limit. The first shows that (noiseless) non-uniform generatability is sufficient for noisy generatability in the limit.
    Theorem (Informal).
    If a class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is (noiseless) non-uniformly generatable, then it is noisily generatable in the limit.
    Since Li et al. (2024) prove that all countable classes are (noiseless) non-uniform generatable, this result also shows that all countable classes are noisily generatable in the limit. In addition, we also give a sufficiency condition in terms in terms of a stronger notion of noisy generatability called “uniform noisy generatability.”
    Theorem (Informal).
    If there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
    One can think of this latter sufficiency condition as the analog of Theorem 3.10 in Li et al. (2024) which shows that the ability to write a class as the finite union of (noiseless) uniformly generatable classes is sufficient for (noiseless) generatability in the limit.
While our theorem statements look similar to that of Li et al. (2024), our proof techniques are different due to the fact that the Noisy Closure dimension is scale-sensitive. This difference manifests even when characterizing noisy uniform generation. In particular, our uniform generator effectively requires combining different generators for each noise-level whereas the noiseless uniform generator from Li et al. (2024) does not.
1.1 Related Work
Language Identification in the Limit. In his seminal 1967 paper, E Mark Gold introduced the model of “Language Identification in the Limit.” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subset U$ for all $i\in\mathbbm{N}$. An adversary picks a language $K\in C$, and begins to enumerate the strings in $K$ one by the one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses an index $i_{t}\in\mathbb{N}$ with the hope that $L_{i_{t}}=K$. The player has identified $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $L_{i_{s}}=K$. The class $C$ is identifiable in the limit, if the player can identify all $K\in C.$ Gold showed that while all finite language families are identifiable in the limit, there are simple countable language families which are not. In a follow-up work, Angluin (1979, 1980) provides a precise characterization of language families $C$ for which language identification in the limit possible. The results by Gold and Angluin emphasized the impossibility of language identification in the limit by ruling out the vast majority of language families. Since Gold’s seminal work on language identification in the limit, there has been extensive follow-up work on this model and we refer the reader to the excellent survey by Lange et al. (2008).
Language Identification from Noisy Examples. In addition to the work by Angluin (1979, 1980), a different line of follow-up work focused on extending Gold’s model to account for noisy example streams. There have been several proposed noise models, and we review a few of them here. The most standard noise model for language identification in the limit allows the adversary to first pick an enumeration of its chosen language $K\in L$ and then insert a finite number of strings that do not belong to $K$ (Baliga et al., 1992; Fulk and Jain, 1989; Schäfer, 1985; Case et al., 1997; Stephan, 1997; Lange and Grieser, 2002). We use the term insert to emphasize that every $w\in K$ must still be present in the sequence chosen by the adversary. Such noisy streams are often referred to as noisy texts. Jain (1994) go beyond the finite nature of the noise by considering sequences of strings with an infinite number of noisy incorrect strings, where the amount of noise is measured using certain density notions. In a different direction, Mukouchi and Sato (2003) and Tantini et al. (2006) model noise at the string-level by defining a distance metric on the space of all strings and allowing the adversaries to insert and delete strings in the original enumeration of $K$ which are at most some distance away from some true string in $K$.
Language Generation in the Limit. Inspired by large language models and recent interest in generative machine learning, Kleinberg and Mullainathan (2024) study the problem of language generation in the limit. In this problem, the adversary also picks a language $K\in C$, and begins to enumerate the strings one by the one to the player in rounds $t=1,2,\dots$. However, now, after observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses a string $\hat{w}_{t}\in U$ with the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. Kleinberg and Mullainathan (2024) prove a strikingly different result than Gold and Angluin – generation in the limit is possible for every countable language family $C$. This positive result has spurred a number of follow-up works, which we briefly review below.
Kalavasis et al. (2024b) study generation in the stochastic setting, where the positive examples revealed to the generator are sampled i.i.d. from some unknown distribution. In this model, they study the trade-offs between generating with breadth and generating with consistency and show that in general, achieving both is impossible, resolving an open question posed by Kleinberg and Mullainathan (2024) for a large family of language models. More recently, Charikar and Pabbaraju (2024) and Kalavasis et al. (2024a) further formalize this tension between consistency and breadth by defining various notions of breadth and providing complete characterizations of which language families are generatable in the limit with breadth.
In a different direction, and most closely related to our work, Li et al. (2024) reinterpret the results of Kleinberg and Mullainathan (2024) through a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over a countable example space $\mathcal{X}$. By doing so, Li et al. (2024) extend the results of Kleinberg and Mullainathan (2024) beyond language generation while also formalizing two stronger settings of generation they term “uniform” and “non-uniform” generation. Unlike Kleinberg and Mullainathan (2024) and Kalavasis et al. (2024a) who focus on computable learners and countable language families, Li et al. (2024) place no such restrictions and provide complete, information-theoretic characterizations of which hypothesis classes are uniformly and non-uniformly generatable.
2 Preliminaries
2.1 Notation
Let $\mathcal{X}$ denote a countable example space and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ denote a binary hypothesis class. Let $\mathcal{X}^{\star}$ denote the set of all finite subsets of $\mathcal{X}$. For any $h\in\mathcal{H}$, an enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that $\bigcup_{i\in\mathbbm{N}}\{x_{i}\}=\operatorname{supp}(h)$. In other words, for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbbm{N}$ such that $x_{i}=x$ and for every $i\in\mathbb{N}$, we have that $x_{i}\in\operatorname{supp}(h).$ For any $h\in\mathcal{H}$, a noisy enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbb{N}$ such that $x_{i}=x$ and
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\not\in\operatorname{supp}(h)\}<\infty.$
For a finite sequence of examples $x_{1},\dots,x_{d}$ and $n\in\mathbb{N}\cup\{0\}$, define
$\mathcal{H}(x_{1},\dots,x_{d},n):=\{h\in\mathcal{H}:|\{x_{1},\dots,x_{d}\}\cap\operatorname{supp}(h)|\geq d-n\}.$
For any class $\mathcal{H}$, define its induced closure operator at noise level $n$ as $\langle\cdot\rangle_{\mathcal{H},n}$ such that
$\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}:=\begin{cases}\bigcap_{h\in\mathcal{H}(x_{1},\dots,x_{d},n)}\operatorname{supp}(h),&\text{if $|\mathcal{H}(x_{1},\dots,x_{d},n)|\geq 1$}\\ \bot,&\text{if $|\mathcal{H}(x_{1},\dots,x_{d},n)|=0$}\end{cases}.$
Note that $\langle\cdot\rangle_{\mathcal{H},0}$ is exactly the closure operator from Section 2.1 of Li et al. (2024). We will make the following assumption about hypothesis classes.
Assumption 1 (Uniformly Unbounded Support (UUS)).
A hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfies the Uniformly Unbounded Support (UUS) property if $|\operatorname{supp}(h)|=\infty$ for every $h\in\mathcal{H}$.
Remark 2.1.
The UUS assumption is mainly needed for bookkeeping purposes to prevent the adversary from exhausting all positive examples and thus making the task of generating unseen positive examples impossible. This is consistent with the assumption that each language is countable infinite in size in the work of Kleinberg and Mullainathan (2024).
Finally, we will let $[N]:=\{1,\dots,N\}$ and abbreviate a finite sequence $x_{1},\dots,x_{n}$ as $x_{1:n}.$
2.2 Generatability
In this paper, we adopt the learning-theoretic framework for generation introduced by Li et al. (2024). To that end, we define a generator as a deterministic map from a finite sequence of examples to a new example.
Definition 2.1 (Generator).
A generator is a map $\mathcal{G}:\mathcal{X}^{\star}\rightarrow\mathcal{X}$ that takes a finite sequence of examples $x_{1},x_{2},\dots$ and outputs a new example $x$.
2.2.1 Noiseless Generation
In the noiseless setting, an adversary plays a game against a generator $\mathcal{G}$. Before the game begins, the adversary picks a hypothesis $h\in\mathcal{H}$ and a sequence of examples $x_{1},x_{2},\dots$ such that $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h).$ The adversary reveals the examples of stream to $\mathcal{G}$ one at a time, and the goal of the generator is to eventually output new, unseen positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$
Depending on how one quantifies “eventually,” one can get various notions of noiseless generatability. For example, if one requires to generator to perfectly generate new unseen example after observing $d\in\mathbbm{N}$ examples regardless of the hypothesis or stream chosen by the adversary, this is called as “uniform generation.”
Definition 2.2 (Uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
If the number of positive examples required before perfect generation can depend on the hypothesis chosen by the adversary, but not the stream, then this is “non-uniform generation.”
Definition 2.3 (Non-uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly generatable if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists a $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Finally, if the number of positive examples before perfect generation can depend on both the hypothesis and the stream selected by the adversary, this is called “generation in the limit.”
Definition 2.4 (Generatability in the Limit (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is generatable in the limit if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, and any enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists a $t^{\star}\in\mathbbm{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Note that uniform generation implies non-uniform generation which implies generation in the limit.
2.2.2 Noisy Generation
In the noisy model, we consider the following game. Like before, the adversary hypothesis $h\in\mathcal{H}$, and a sequence of positive examples $z_{1},z_{2},\dots\subseteq\operatorname{supp}(h)$. But now, the picks a noise level $n^{\star}\in\mathbb{N}$, and inserts at most $n^{\star}$ negative examples in $z_{1},z_{2},\dots$ to obtain a noisy stream $x_{1},x_{2},\dots$. The adversary then presents the examples in the noisy stream to generator $\mathcal{G}$ one at a time. Without knowledge of the noise-level or the location of the negative examples, the goal of the generator is to eventually output new, positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$.
Like the noiseless case, there are several definitions for noisy generatability based on how one quantifies “eventually.” The most extreme case is what we term “uniform noisy generatability.”
Definition 2.5 (Uniform Noisy Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noisily generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}\cap\operatorname{supp}(h)|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Here, the generator must perfectly generate after seeing $d\in\mathbb{N}$ positive example, where $d$ can only depend on the class $\mathcal{H}$ itself, and thus is uniform over the noise-level $n^{\star}$, the hypothesis $h\in\mathcal{H}$, and the stream $x_{1},x_{2},\dots$ selected by the adversary. In general, as we show in Section 3.1, uniform noisy generatability is impossible even for simple classes with just two hypotheses. This motivates weakening the definition by allowing $d$ to depend on the noise-level. We call this setting “Noisy Uniform Generatability.”
Definition 2.6 (Noisy Uniform Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily uniformly generatable, if there exists a generator $\mathcal{G}$ such that for every noise level $n^{\star}\in\mathbb{N}$, there exists a $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Noisy uniform generatability does not suffer from the same hardness of uniform noisy generatability. In fact, in Section 3.2, we show that all finite classes are uniformly noisily generatable. Nevertheless, we can continue weakening the notion of noisy generatability, by allowing $d$ to also depend on the hypothesis chosen by the adversary.
Definition 2.7 (Noisy Non-uniform Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily non-uniformly generatable, if there exists a generator $\mathcal{G}$, such that for every noise level $n^{\star}\in\mathbb{N}$ and any $h\in\mathcal{H}$ there exists $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
The term “non-uniform” here is used to refer to the fact that the number of unique examples needed by the generator before perfect generation can depend on the selected $h\in\mathcal{H}$, and hence it is “non-uniform” over the hypothesis class, unlike the previous two definitions. Finally, the weakest from of noisy generation allows $d$ to depend on the noise-level, the hypothesis, and the stream selected by the adversary.
Definition 2.8 (Noisy Generatability in the Limit).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily generatable in the limit, if there exists a generator $\mathcal{G}$, such that for every $h\in\mathcal{H}$ and any noisy enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists $t^{\star}\in\mathbbm{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Note that in Definition 2.8, we require the noisy stream picked by the adversary to still contain every example in the support of the selected hypothesis. This is consistent with the model of “noisy texts” from the literature in language identification where one allows the adversary to insert noisy examples in the stream, as opposed to replacing positive examples (Stephan, 1997).
Remark 2.2.
We highlight that in all of these definitions, the generator is not aware of the noise-level or the noisy examples picked by the adversary. However, for all but “Uniform Noisy Generatability,” we allow the number of positive examples needed before perfect generation to depend on the noise-level in the stream.
Remark 2.3.
The astute reader might notice that there is actually a fifth setting of noisy generatability that we did not cover. In this fifth setting, which we will call “Non-uniform Noisy Generatability,” the number of positive examples needed before perfect generation can depend on the hypothesis chosen by the adversary, but must still be uniform over the noise-level and the noisy example stream chosen by the adversary. However, as in the case of uniform noisy generatability, non-uniform noisy generatability is still too strong as there exists a class of just two hypotheses that is not non-uniformly noisily generatable. We refer the reader to Appendix A for more details.
2.3 Summary of Results for Noiseless Generation
In the noiseless setting, Kleinberg and Mullainathan (2024) initiated the study of generation in the limit by showing that all countable classes are generatable in the limit and that all finite classes are uniformly generatable. Following this result, Li et al. (2024) provided a complete characterization of uniform generation in terms of a new combinatorial parameter termed the Closure dimension.
Definition 2.9 (Closure dimension (Li et al., 2024)).
The Closure dimension of $\mathcal{H}$, denoted $\operatorname{C}(\mathcal{H})$, is the largest natural number $d\in\mathbbm{N}$ for which there exists distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}|<\infty.$ If this is true for arbitrarily large $d\in\mathbbm{N}$, then we say that $\operatorname{C}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{C}(\mathcal{H})=0.$
In particular, a class is uniformly generatable if and only its its Closure dimension is finite.
Theorem 2.1 (Theorem 3.3. in Li et al. (2024)).
A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is uniformly generatable if and only if $\operatorname{C}(\mathcal{H})<\infty$.
In addition, Li et al. (2024) defined an intermediate setting termed non-uniform generatability, and provided a complete characterization of which classes are non-uniformly generatable.
Theorem 2.2 (Theorem 3.5 in Li et al. (2024)).
A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is non-uniformly generatable if and only if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{n}$ and $\operatorname{C}(\mathcal{H}_{n})<\infty$ for every $n\in\mathbb{N}$.
As a corollary, Li et al. (2024) establish that all countable classes are actually non-uniformly generatable. Charikar and Pabbaraju (2024) also established this result independently. With regards to generatability in the limit, Kleinberg and Mullainathan (2024) prove that all countable classes are generatable in the limit. Li et al. (2024) provide an alternate sufficiency condition in terms of the Closure dimension.
Theorem 2.3 (Theorem 3.10 in Li et al. (2024)).
A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is generatable in the limit if there exists a finite sequence of classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{n}$ such that $\mathcal{H}=\bigcup_{i=1}^{n}\mathcal{H}_{i}$ and $\operatorname{C}(\mathcal{H}_{i})<\infty$ for all $i\in[n].$
In this paper, we provide analogous results for noisy generation in terms of a different combinatorial parameter termed the Noisy Closure dimension.
3 Towards Characterizations of Noisy Generation
3.1 Uniform Noisy Generatability
We start by providing a characterization of the harshest form of noisy generation – uniform noisy generatability.
Theorem 3.1 (Characterization of Uniform Noisy Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. The following statements are equivalent.
  • (i)
    $\mathcal{H}$ is uniformly noisily generatable.
  • (ii)
    $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=\infty.$
Theorem 3.1 is a hardness result – it shows that even finite classes with just two hypotheses may not be uniformly noisily generatable. In fact, one way to interpret Theorem 3.1 is that uniform noisy generation is only possible for trivial classes where the generator can perfectly generate without observing any examples from the adversary. Indeed, this is what condition (ii) in Theorem 3.1 implies, since if condition (ii) is true, the generator can simply compute $\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ and always play from this set.
Proof.
(of sufficiency in Theorem 3.1) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose that $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=\infty.$ Then, the generator $\mathcal{G}$ which simply computes $S=\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ and plays arbitrarily from $S\setminus\{x_{1},\dots,x_{t}\}$ on every round $t\in\mathbb{N}$ generates perfectly (according to Definition A.1) from round $t=1$ itself, completing the proof. ∎
Proof.
(of necessity in Theorem 3.1) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose that $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}<\infty.$ We need to show that $\mathcal{H}$ is not uniformly noisily generatable. Let $\mathcal{F}\subseteq\mathcal{H}$ be the smallest subset of $\mathcal{H}$ such that
$\Bigl{|}\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\Bigl{|}<\infty.$
Then, by definition, we know that for every $f\in\mathcal{F}$, we have that
$\Bigl{|}\bigcap_{h\in\mathcal{F}\setminus f}\operatorname{supp}(h)\Bigl{|}=\infty.$
We will show that $\mathcal{F}$ is not uniformly noisily generatable which will imply that $\mathcal{H}$ is not uniformly noisily generatable. To that end, we need to show that for every $\mathcal{G}$ and any $d\in\mathbb{N}$, there exists a $h\in\mathcal{F}$ and a sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$
such that for every $t\in\mathbb{N}$ where |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|=d$, there exists an $s\geq t$ such that $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$. To that end, fix a generator $\mathcal{G}$ and a number $d\in\mathbbm{N}$. Fix some $f\in\mathcal{F}$ and let $x_{1},\dots,x_{d}$ be a sequence of $d$ unique examples in $\bigcap_{h\in\mathcal{F}\setminus f}\operatorname{supp}(h)$. Let $z_{1},\dots,z_{d}$ be any $d$ unique examples in $\operatorname{supp}(f)\setminus\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)$ and consider the sequence of $2d$ unique examples obtained by concatenation $z_{1},\dots,z_{d}$ to the end of $x_{1},\dots,x_{d}$. Denote this new sequence by $x_{1:2d}.$ Let $q=|\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)|$ and let $v_{1},\dots,v_{q}$ be its elements sorted in its natural order. Concatenate $v_{1},\dots,v_{q}$ to the end of $x_{1:2d}$ and denote this new sequence of unique examples as $x_{1:2d+q}.$ Observe that for every $h\in\mathcal{F}$, we have that $|\operatorname{supp}(h)\cap\{x_{1},\dots x_{2d+q}\}|\geq d.$ Let $\hat{x}_{2d+q}=\mathcal{G}(x_{1},\dots,x_{2d+q})$ and suppose without loss of generality that $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}.$ Let $h^{\star}\in\mathcal{F}$ be a be hypothesis such that $\hat{x}_{2d+q}\notin\operatorname{supp}(h^{\star}).$ Such a hypothesis must exist because $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}$ and $\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\subseteq\{x_{1},\dots,x_{2d+q}\}.$ Let $x_{2d+q+1},x_{2d+q+2},\dots$ be any completion of the stream such that $\{x_{2d+q+t}\}_{t=1}^{\infty}\subseteq\operatorname{supp}(h^{\star})$ and $\{x_{2d+q+t}\}_{t=1}^{\infty}\cap\{x_{t}\}_{t=1}^{2d+q}=\emptyset.$
We are now ready to complete the proof. Let $h^{\star}$ and $x_{1},x_{2},\dots$ be the hypothesis and sequence chosen above. Then, by definition, observe that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}=\sum_{t=1}^{2d+q}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}\leq d<\infty.$
Moreover, we also know that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots x_{2d+q}\}|\geq d$ and the first $2d+q$ examples are distinct, implying that there exists exactly one time point $t\leq 2d+q$ such that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots x_{2d+q}\}|=d.$ Finally, noting that $2d+q\geq t$ and that $\mathcal{G}(x_{1},\dots,x_{2d+q})\notin\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{2d+1}\}$ completes the proof that $\mathcal{F}$ is not uniformly noisily generatable since $\mathcal{G}$ and $d$ were chosen arbitrarily. ∎
3.2 Noisy Uniform Generatability
Theorem 3.1 shows that obtaining guarantees that are uniform over the noise level is generally hopeless. In this section, we study the easier setting of noisy uniform generation, where the number of examples needed for perfect generation can depend on the noise level. At a high-level the results in this section extend the techniques from Li et al. (2024) to account for noisy example streams. As such, we first define a noisy analog of the Closure dimension.
Definition 3.1 ($n$-Noisy Closure dimension).
The Noisy Closure dimension of $\mathcal{H}$ at level $n\in\mathbbm{N}$, denoted $\operatorname{NC}_{n}(\mathcal{H})$, is the largest natural number $d$ for which there exists distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}|<\infty.$ If this is true for arbitrarily large $d\in\mathbbm{N}$, then we say that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{NC}_{n}(\mathcal{H})=0.$
Unlike the Closure dimension, the Noisy Closure dimension is scale-sensitive, it is defined with respect to every noise-level $n\in\mathbb{N}$. Scale-sensitive dimensions are not new to learning theory, but have also been defined and used to characterize other properties of hypothesis classes like PAC and online learnability for regression problems (Rakhlin et al., 2015; Bartlett et al., 1994). Our main theorem in this section uses the Noisy Closure dimension to provide a complete characterization of which classes are noisily uniform generatable.
Theorem 3.2 (Characterization of Noisy Uniform Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. The following statements are equivalent.
  • (i)
    $\mathcal{H}$ is noisily uniformly generatable.
  • (ii)
    $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$
We start with the necessity direction.
Proof.
(of necessity in Theorem 3.2) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose there exists a $n\in\mathbb{N}$ such that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ We need to show that this means that $\mathcal{H}$ is not noisily uniformly generatable. In particular, we need to show that for every generator $\mathcal{G}$, there exists a noise level $n^{\star}\in\mathbb{N}$ such that for every $d\in\mathbbm{N}$, there exists a hypothesis $h\in\mathcal{H}$ and a sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
so that for every $t\in\mathbbm{N}$ with |$|\{x_{1},\dots,x_{t}\}|=d$, there exists $s\geq t$ where $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$.To that end, let $\mathcal{G}$ be any generator and consider the noise level $n^{\star}=n$ so that $\operatorname{NC}_{n^{\star}}(\mathcal{H})=\infty.$ Fix any $d\in\mathbb{N}.$ We will now pick such a hypothesis $h^{\star}\in\mathcal{H}$ and a valid stream $x_{1},x_{2},\dots.$
Since $\operatorname{NC}_{n^{\star}}(\mathcal{H})=\infty$, we know there exists a $d^{\star}\geq d$ and a sequence of distinct examples $x_{1},\dots,x_{d^{\star}}$ such that $\langle x_{1},\dots,x_{d^{\star}}\rangle_{\mathcal{H},n^{\star}}\neq\bot$ and $|\langle x_{1},\dots,x_{d^{\star}}\rangle_{\mathcal{H},n^{\star}}|<\infty.$ Let $q=|\langle x_{1},\dots,x_{d^{\star}}\rangle_{\mathcal{H},n^{\star}}\setminus\{x_{1},\dots,x_{d^{\star}}\}|$ and $z_{1},\dots,z_{q}$ be the elements of $\langle x_{1},\dots,x_{d^{\star}}\rangle_{\mathcal{H},n^{\star}}\setminus\{x_{1},\dots,x_{d^{\star}}\}$ sorted in their natural order. Let $x_{1:d^{\star}+q}$ denote the sequence obtained by concatenating $z_{1},\dots,z_{q}$ to the end of $x_{1},\dots,x_{d^{\star}}.$ Let $\hat{x}_{d^{\star}+q}=\mathcal{G}(x_{1:d^{\star}+q})$ and suppose without loss of generality that $\hat{x}_{d^{\star}+q}\notin\{x_{1},\dots,x_{d^{\star}+q}\}.$ Let $h^{\star}\in\mathcal{H}(x_{1:d^{\star}},n^{\star})$ be any hypothesis such that $\hat{x}_{d^{\star}+q}\notin\operatorname{supp}(h^{\star}).$ Such a hypothesis must exist because $\hat{x}_{d^{\star}+q}\notin\{x_{1},\dots,x_{d^{\star}+q}\}$ and $\langle x_{1},\dots,x_{d^{\star}}\rangle_{\mathcal{H},n^{\star}}\subseteq\{x_{1},\dots,x_{d^{\star}+q}\}.$ Finally, let $x_{d^{\star}+q+1},x_{d^{\star}+q+2},\dots$ be any completion of the stream such that $\{x_{d^{\star}+q+t}\}_{t=1}^{\infty}\subseteq\operatorname{supp}(h^{\star})$ and $\{x_{d^{\star}+q+t}\}_{t=1}^{\infty}\cap\{x_{t}\}_{t=1}^{d^{\star}+q}=\emptyset.$
We now complete the proof. Let $h^{\star}$ and $x_{1},x_{2},\dots$ be the hypothesis and stream chosen above. First, observe that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}=\sum_{t=1}^{d^{\star}}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}\leq n^{\star}.$
Second, there exists only one $t\in\mathbb{N}$, namely $t=d$, such that |$|\{x_{1},\dots,x_{t}\}|=d$ as the first $d^{\star}+1>d$ examples of the stream are distinct by construction. However, observe that by construction, at time point $s=d^{\star}+q\geq d$, we have that $\mathcal{G}(x_{1},\dots,x_{s})=\hat{x}_{d^{\star}+q}\notin\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{s}\}$ by choice of $h^{\star}.$ Since $\mathcal{G}$ and $d\in\mathbb{N}$ were picked arbitrarily, this is true for all $\mathcal{G}$ and $d\in\mathbb{N}$, which completes the proof. ∎
Finally, we prove the sufficiency direction.
Proof.
(of sufficiency in Theorem 3.2) The key intuition is that the generator, without knowing the adversary’s noise level, continually computes the largest noise level for which it has observed enough unique examples to generate perfectly. Eventually, its calculated noise level will surpass and thus account for the adversary’s noise level, ensuring that the generator is perfect from then on.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. We need to show that if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N},$ then $\mathcal{H}$ is noisily uniformly generatable. For a sequence $x_{1},x_{2},\dots$, let $d_{t}:=|\{x_{1},\dots,x_{t}\}|$ denote the number of distinct examples amongst the first $t$ examples. Consider the following generator $\mathcal{G}.$ At time $t\in\mathbb{N}$, $\mathcal{G}$ computes
$n_{t}:=\max\{n\in\mathbb{N}:\operatorname{NC}_{n}(\mathcal{H})<d_{t}\}\cup\{0\}.$
If $n_{t}=0,$ $\mathcal{G}$ plays any $x\in\mathcal{X}.$ Otherwise, it plays from $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n_{t}}\setminus\{x_{1},x_{2},\dots,x_{t}\}.$ We now show that $\mathcal{G}$ is a noisy uniform generator. To that end, let $n^{\star}\in\mathbb{N}$ be the noise level chosen by the adversary. We claim that $\mathcal{G}$ can perfectly generate after observing $\operatorname{NC}_{n^{\star}}(\mathcal{H})+1$ distinct examples. To see why, let $h\in\mathcal{H}$ be any hypothesis, and $x_{1},x_{2},\dots$ be any stream such that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star}.$
Without loss of generality, suppose there exists $t^{\star}\in\mathbb{N}$ such that $d_{t^{\star}}=\operatorname{NC}_{n^{\star}}(\mathcal{H})+1.$ Fix any $s\geq t^{\star}$. By definition of $n_{t}$, it must be the case that $n_{s}\geq n^{\star}.$ Thus, $h\in\mathcal{H}(x_{1},\dots,x_{s},n_{s})$ and $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}\subseteq\operatorname{supp}(h)$. Moreover, because $d_{s}>\operatorname{NC}_{n_{s}}(\mathcal{H})$, again by definition, it must be the case that $|\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}|=\infty.$ Accordingly, $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}\setminus\{x_{1},\dots,x_{s}\}\neq\emptyset$ and $\mathcal{G}$ is guaranteed to output an example in $\operatorname{supp}(h)$ on round $s$. Since $s\geq t^{\star}$ was chosen arbitrarily, this is true for all such $s$, completing the proof. ∎
Unlike uniform noisy generation, our next result shows that all finite classes are noisily uniformly generatable.
Corollary 3.3 (All Finite Classes are Noisily Uniformly Generatable).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is finite, then $\mathcal{H}$ is noisily uniformly generatable.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be a finite hypothesis class with $q:=|\mathcal{H}|$ satisfying the $\operatorname{UUS}$ property. To show that $\mathcal{H}$ is noisily uniformly generatable, it suffices to show that $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$ For any subset $V\subseteq\mathcal{H},$ define
$\langle\emptyset\rangle_{V}=\bigcap_{h\in V}\operatorname{supp}(h)\text{ and }\mathcal{F}=\{V\subseteq\mathcal{H}:|\langle\emptyset\rangle_{V}|<\infty\}$
to be the set of examples common to all hypotheses in $V$ and the subsets of $\mathcal{H}$ whose intersection of supports has finite cardinality, respectively. Then, let
$d:=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|,$
be the maximum size of a finite set of examples common to a subset of $\mathcal{H}$. Note that $d$ is finite because $\mathcal{H}$ is finite. Fix any noise level $n\in\mathbb{N}$ and for the sake of contradiction, suppose $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ This means that for every $s\in\mathbb{N}$, there exists $t\geq s$ and a sequence of distinct examples ${x_{1},\dots,x_{t}}$ such that $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|<\infty$. Pick $s=nq+d+1$ and consider the stream $x_{1},\dots,x_{t}$ such that $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}<\infty$ for $t\geq s.$ Recall that $\mathcal{H}(x_{1:t},n)=\{h\in\mathcal{H}:|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|\geq t-n\}$ That is, $\mathcal{H}(x_{1:t},n)$ contains all hypotheses in $\mathcal{H}$ that are inconsistent with $x_{1:t}$ on at most $n$ examples. Since $|\mathcal{H}(x_{1:t},n)|\leq|\mathcal{H}|=q,$ there are at least
$t-nq\geq(nq+d+1)-nq=d+1$
distinct examples in $x_{1:t}$ contained in the support of all hypotheses in $\mathcal{H}(x_{1:t},n)$. In other words, $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|=\langle\emptyset\rangle_{\mathcal{H}(x_{1:t},n)}\geq d+1>d.$ This is a contradiction as $d=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|$ by definition and $\mathcal{H}(x_{1:t},n)\in\mathcal{F}.$ Thus, $\operatorname{NC}_{n}(\mathcal{H})<nq+d+1<\infty$ for all $n\in\mathbb{N},$ completing the proof. ∎
We end this section by establishing that noisy uniform generatability is strictly harder than noiseless uniform generatability. This is unlike PAC and online learnability for classification problems, where typically the noise-free (realizable) and noisy (agnostic) characterizations of learnability are equivalent (Littlestone, 1987; Ben-David et al., 2009).
Lemma 3.4 (Uniform Generatability $\neq$ Noisy Uniform Generatability).
Let $\mathcal{X}$ be countable. There exists a countable class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfying the UUS property such that $\operatorname{C}(\mathcal{H})=0$ but $\operatorname{NC}_{1}(\mathcal{H})=\infty.$
Proof.
Let =$\mathbb{P}=\{p_{1},p_{2},\dots\}$ be the set of primes, $S_{q}=\{-q^{n}:n\in\mathbb{N}\}$, $\mathbb{S}=\bigcup_{q\in\mathbb{P}}S_{q}$ and $\mathcal{X}=\mathbb{P}\cup\mathbb{S}$. Let the support of each hypothesis $h_{p}$ be
$\operatorname{supp}(h_{p})=\mathbb{P}\setminus\{p\}\cup\mathbb{S}\setminus S_{p}$
and :=$\mathcal{H}:=\{h_{p}:p\in\mathbb{P}\}.$ Observe that $\mathcal{H}$ satisfies the $\operatorname{UUS}$ property and $\mathcal{X}$ is countable.
We now prove that $\operatorname{C}(\mathcal{H})=0.$ Suppose the first element in the sequence is $x_{1}=-p^{n}$ for some $p\in\mathbb{P}$ and $n\in\mathbb{N}.$ Then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{m}:m\in\mathbb{N}\ \}|=\infty.$ If the first element in the sequence is $x_{1}=p\in\mathbb{P},$ then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{n}:n\in\mathbb{N}\ \}|=\infty.$ Note that these are the only two possible cases as it must be the case that $x_{1}\in\operatorname{supp}(h)$ for some $h\in\mathcal{H}.$
Now, we show that $\operatorname{NC}_{1}(\mathcal{H})=\infty.$ Fix some $d\in\mathbb{N}.$ Consider the stream $p_{1},p_{2},\dots,p_{d}$ of the first $d$ prime numbers. Then, $|\langle p_{1},\dots,p_{d}\rangle_{\mathcal{H},1}|=|\emptyset|=0,$ due to the fact that $\mathcal{H}(p_{1},\dots,p_{d},1)=\mathcal{H}$ and the intersection of the supports of all $h\in\mathcal{H}$ is empty. So, $\operatorname{NC}_{1}(\mathcal{H})\geq d.$ Since $d\in\mathbb{N}$ was picked arbitrarily, this is true for all $d\in\mathbbm{N}$ and therefore $\mathcal{H}$ is not noisily uniformly generatable. ∎
Lemma 3.4 shows a strong separation between noise-free and noisy uniform generation – there is a class that is uniformly generatable, but not noisily uniform generatable even when the adversary is only allowed to perturb at most one example.
3.3 Noisy Non-uniform Generatability
Similar to the characterization of non-uniform generatability in the noise-free setting, we can use the noisy uniform generatability to provide sufficiency and necessary conditions for noisy non-uniform generatability.
Lemma 3.5 (Sufficiency for Noisy Non-uniform Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}$, then $\mathcal{H}$ is noisily non-uniformly generatable.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any class satisfying the $\operatorname{UUS}$ property. Suppose there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}$. For a sequence $x_{1},x_{2},\dots$, let $d_{t}:=|\{x_{1},\dots,x_{t}\}|$ denote the number of distinct examples amongst the first $t$ examples. Consider the following generator $\mathcal{G}.$ At time $t\in\mathbb{N}$, $\mathcal{G}$ computes
$j_{t}=\max\{i\in\mathbb{N}:\operatorname{NC}_{i}(\mathcal{H}_{i})<d_{t}\}\cup\{0\}.$
If $j_{t}=0,$ $\mathcal{G}$ plays any $x\in\mathcal{X}.$ Otherwise, it plays from $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H}_{j_{t}},j_{t}}\setminus\{x_{1},x_{2},\dots,x_{t}\}.$ We claim that $\mathcal{G}$ is a noisy non-uniform generator for $\mathcal{H}$. To see why, let $n^{\star}\in\mathbb{N}$ and $h\in\mathcal{H}$ be the noise level and hypothesis picked by the adversary. Let $i^{\star}\in\mathbb{N}$ be the smallest index such that $h\in\mathcal{H}_{i^{\star}}$ and define $j^{\star}=\max\{i^{\star},n^{\star}\}$. We claim that $\mathcal{G}$ perfectly generates after observing $\operatorname{NC}_{j^{\star}}(\mathcal{H}_{j^{\star}})+1$ distinct examples. To that end, let $x_{1},x_{2},\dots$ be any sequence such that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star}.$
Without loss of generality, suppose there exists $t^{\star}\in\mathbb{N}$ such that $d_{t^{\star}}=\operatorname{NC}_{j^{\star}}(\mathcal{H}_{j^{\star}})+1.$ Fix any $s\geq t^{\star}$. By definition of $j_{t}$, it must be the case that $j_{s}\geq j^{\star}\geq n^{\star}.$ Moreover, because $j_{s}\geq i^{\star}$, we also have that $h\in\mathcal{H}_{j_{s}}$. Together, this means that $h\in\mathcal{H}_{j_{s}}(x_{1},\dots,x_{s},j_{s})$ and $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}\subseteq\operatorname{supp}(h)$. Also, because $d_{s}>\operatorname{NC}_{j_{s}}(\mathcal{H}_{j_{s}})$ it must be the case that $|\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}|=\infty.$ Accordingly, $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}\setminus\{x_{1},\dots,x_{s}\}\neq\emptyset$ and $\mathcal{G}$ is guaranteed to output an example in $\operatorname{supp}(h)$ on round $s$. Since $s\geq t^{\star}$ was chosen arbitrarily, this is true for all such $s$, completing the proof. ∎
As a corollary of Lemma 3.5 and Corollary 3.3, we can show that all countable classes are noisily non-uniformly generatable, and hence noisily generatable in the limit. In some sense, this result shows that amongst countable classes, noisy generation comes for free!
Corollary 3.6 (All Countable Classes are Noisily Non-uniformly Generatable).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is countably infinite, then $\mathcal{H}$ is noisily non-uniformly generatable, and therefore also noisily generatable in the limit.
Proof.
Since noisy non-uniform generatability implies noisy generatability in the limit, it suffices to show that every countable $\mathcal{H}$ is noisily non-uniformly generatable. Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any countable class satisfying the UUS property. Let $h_{1},h_{2},\dots$ be some fixed enumeration of $\mathcal{H}$ and define $\mathcal{H}_{n}=\{h_{i}:i\leq n\}$ for all $n\in\mathbb{N}$. Then, observe that $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ and that $\mathcal{H}=\bigcup_{n=1}^{\infty}\mathcal{H}_{n}$. Next, observe that $|\mathcal{H}_{n}|=n<\infty$, which, using Theorem 3.2, implies that $\operatorname{NC}_{n}(\mathcal{H}_{n})<\infty$. Finally, Lemma 3.5 gives that $\mathcal{H}$ is noisily non-uniformly generatable. ∎
Corollary 3.6 with Lemma 3.4 also establishes that noisy uniform generatability is strictly harder than noisy non-uniform generatability. We now move to our necessity condition for noisy non-uniform generatability.
Lemma 3.7 (Necessity for Noisy Non-uniform Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is noisily non-uniformly generatable, then for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any noisily non-uniformly generatable class satisfying the $\operatorname{UUS}$ property. Suppose that $\mathcal{H}$ is not uniformly noisily generatable and $\mathcal{G}$ be a noisy non-uniform generator for $\mathcal{H}$. Fix $n\in\mathbb{N}$. For every $h\in\mathcal{H}$, let $d_{h,n}\in\mathbb{N}$ be the smallest natural number such that for any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n,$ if there exists a $t\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t}\}|=d_{h,n}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t$. Let $\mathcal{H}_{i}=\{h\in\mathcal{H}:d_{h,n}\leq i\}$ for all $i\in\mathbb{N}$. Note that $\mathcal{H}=\bigcup_{i\in\mathbb{N}}\mathcal{H}_{i}$ because $d_{h,n}<\infty$ for all $h\in\mathcal{H}.$ Moreover, we have that $\mathcal{H}_{i}\subseteq\mathcal{H}_{i+1}$ for all $i\in\mathbb{N}.$ Finally, for every $i\in\mathbb{N}$, observe that $\mathcal{G}$ is a uniform generator for $\mathcal{H}_{i}$ at noise level $n\in\mathbb{N}$, implying that $\operatorname{NC}_{n}(\mathcal{H}_{i})<i<\infty.$
Note the sufficiency and necessary conditions in Lemmas 3.5 and 3.7 respectively are not matching. Indeed, the condition in Lemma 3.5 is stronger than that in Lemma 3.7. We leave as an open question a complete characterization of noisy non-uniform generatability.
3.4 Noisy Generatability in the Limit
Corollary 3.6 showed that all countable classes are noisily generatable in the limit. Here, we provide alternate sufficiency conditions for noisy generatability in the limit. Our first result shows that (noiseless) non-uniform generatability is sufficient for noisy generatability in the limit. Since all countable classes are (noiseless) non-uniform generatable (Li et al., 2024), this result also shows that all countable classes noisily generatable in the limit.
Theorem 3.8 (Non-uniform Generatability $\implies$ Noisy Generatability in the Limit).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any class satisfying the UUS property. If $\mathcal{H}$ is (noiseless) non-uniformly generatable, then $\mathcal{H}$ is noisily generatable in the limit.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any class satisfying the UUS property. Consider the generator $\mathcal{G},$ which uses a non-uniform generator $\mathcal{Q}$ as in Algorithm 1.
Input: Hypothesis class $\mathcal{H}$ and non-uniform generator $\mathcal{Q}$
1
2for $t=1,2,\dots$ do
3
4      Adversary reveals example $x_{t}$
5      Let $d_{t}=|\{x_{1},\dots,x_{t}\}|$ and $r_{t}\leq t$ be the largest time point such that |$|\{x_{r_{t}},\dots,x_{t}\}|=\lfloor\frac{d_{t}}{2}\rfloor$
6      Initialize $\hat{z}^{t}_{1}=\mathcal{Q}(x_{r_{t}:t})$
7      for $i=1,\dots,r_{t}$ do
8
9            if $\hat{z}^{t}_{i}\notin\{x_{1},\dots,x_{t}\}$ then
10                   $\mathcal{G}$ outputs $\hat{z}^{t}_{i}$ and moves to next round.
11             else
12                   Update $\hat{z}^{t}_{i+1}=\mathcal{Q}(x_{r_{t}},\dots,x_{t},\hat{z}^{t}_{1},\dots,\hat{z}^{t}_{i})$
13
14       end for
15      $\mathcal{G}$ outputs $\hat{z}^{t}_{r_{t}+1}$ and moves to next round.
16 end for
Algorithm 1 Generator $\mathcal{G}$
We will show that $\mathcal{G}$ noisily generates from $\mathcal{H}$ in the limit. Let $h^{\star}\in\mathcal{H}$ and $x_{1},x_{2},\dots$ be the chosen hypothesis and noisy enumeration of $\operatorname{supp}(h^{\star})$ picked by the adversary, respectively. Let $d_{\mathcal{Q}}(h^{\star})\in\mathbbm{N}$ be the number of noiseless distinct elements that $\mathcal{Q}$ needs for $h^{\star}$ to perfectly generate from $\operatorname{supp}(h^{\star})$. Let $t^{\star}\in\mathbb{N}$ be such that for all $t\geq t^{\star}$, we have that $x_{t}\in\operatorname{supp}(h^{\star}).$ Such a $t^{\star}$ must exist because there are most a finite number of noisy examples in the stream. Also, because $x_{1},x_{2},\dots$ is a noisy enumeration, at some point $s^{\star}\in\mathbb{N}$, $r_{s^{\star}}\geq t^{\star}$ and $\lfloor\frac{d_{s^{\star}}}{2}\rfloor\geq d_{\mathcal{Q}}(h^{\star})$. Fix some $s\geq s^{\star}.$ We will prove that $\mathcal{G}$ generates perfectly on round $s$. First, we claim that for all $i\in[r_{s}+1]$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i-1}\}$. Our proof will be by induction. For the base case, consider $i=1$. We need to show that $\hat{z}^{s}_{1}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t}\}$. However, this just follows from the fact that $\hat{z}^{s}_{1}=\mathcal{Q}(x_{r_{t}:t})$, |$|\{x_{r_{t}},\dots,x_{t}\}|\geq d_{\mathcal{Q}}(h^{\star})$, and $\{x_{r_{t}},\dots,x_{t}\}\subseteq\operatorname{supp}(h^{\star}).$ Next, for the induction step, suppose that for all $i\leq m$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h)\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i}\}$. Then, by definition, we have that $\hat{z}_{m+1}^{s}=\mathcal{Q}(x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m})$. The proof of the claim is complete after noting that $\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}\subseteq\operatorname{supp}(h^{\star})$ and |$|\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}|\geq d_{\mathcal{Q}}(h^{\star}).$
Now we will complete the overall proof by showing that the output of $\mathcal{G}$ on round $s$ lies in $\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}$. Suppose that $\mathcal{G}$ outputs $\hat{z}_{j}^{s}$ for some $j\in[r_{t}+1]$. If $j<r_{t}+1$, then it must be the case that Line 7 fired and so by construction $\hat{z}_{j}^{s}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}$. Thus, suppose that $j=r_{t}+1.$ If $j=r_{t}+1$, then it must be the case that $\{\hat{z}_{1}^{s},\dots,\hat{z}_{r_{t}}^{s}\}=\{x_{1},\dots,x_{r_{t}}\}.$ Accordingly, we have that $\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{t}}\}=\{x_{1},\dots,x_{t}\}$, which means that $\hat{z}^{s}_{r_{t}+1}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{t}}\}$ and therefore $\hat{z}^{s}_{r_{t}+1}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}.$ Thus, in either case, $\mathcal{G}$ perfectly generates on round $s$. Since $s\geq s^{\star}$ is arbitrary, our proof is complete. ∎
In similar spirit to Theorem 3.10 from Li et al. (2024), we provide a another sufficiency condition for noisy generatability in the limit in terms of uniform noisy generatability.
Theorem 3.9.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$. By Theorem 3.1, we know that for all $i\in[k]$, we have that
$\Bigl{|}\bigcap_{h\in\mathcal{H}_{i}}\operatorname{supp}(h)\Bigl{|}=\infty.$
For every $i\in[k]$, let $z^{i}_{1},z^{i}_{2},\dots$ denote the elements of $\bigcap_{h\in\mathcal{H}_{i}}\operatorname{supp}(h)$ sorted in their natural ordering. Consider the following generator $\mathcal{G}$. Given any noisy enumeration $x_{1},x_{2},\dots$ and any $t\in\mathbb{N}$, $\mathcal{G}$ first computes
$p_{t}^{i}:=\max\{p\in\mathbb{N}:\{z^{i}_{1},\dots,z^{i}_{p}\}\subseteq\{x_{1},\dots,x_{t}\}\}$
for every $i\in[k]$. Then, $\mathcal{G}$ computes $i_{t}=\operatorname*{arg\,max}_{i\in[k]}p^{i}_{t}.$ Finally, $\mathcal{G}$ plays arbitrarily from $\{z^{i_{t}}_{1},z^{i_{t}}_{2},\dots\}\setminus\{x_{1},\dots x_{t}\}.$ We claim that $\mathcal{G}$ generates from $\mathcal{H}$ in the limit.
To see why, let $h\in\mathcal{H}$ and $x_{1},x_{2},\dots$ be the hypothesis and noisy enumeration selected by the adversary. Let $i^{\star}\in[k]$ be such that $h\in\mathcal{H}_{i^{\star}}$ and $s^{\star}\in\mathbb{N}$ be such that for all $t\geq s^{\star}$, we have that $x_{t}\in\operatorname{supp}(h)$. Such an $s^{\star}$ exists because $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty.$ Let $S^{\star}\subseteq[k]$ be such that $i\in S^{\star}$ if and only if $\{z^{i}_{1},z^{i}_{2},\dots\}\subseteq\{x_{1},x_{2},\dots\}.$ Note that $i^{\star}\in S^{\star}$ by definition of $z^{i^{\star}}_{1},z^{i^{\star}}_{2},\dots$ and the fact that $x_{1},x_{2},\dots$ is a noisy enumeration. By definition of $S^{\star}$ and $p_{t}^{i}$ it must be the case that for all $j\notin S^{\star}$, we have that $\sup_{t\in\mathbb{N}}p_{t}^{j}<\infty.$ Since $k<\infty$, we then get that $\max_{j\notin S^{\star}}\sup_{t\in\mathbb{N}}p_{t}^{j}<\infty.$ On the other hand, for all $i\in S^{\star}$, we have that $p_{t}^{i}\rightarrow\infty.$ More precisely, for every $i\in S^{\star}$, there exists a finite $t_{i}\in\mathbb{N}$ such that $p^{i}_{t_{i}}\geq\max_{j\notin S^{\star}}\sup_{t\in\mathbb{N}}p_{t}^{j}.$ Since $|S^{\star}|<\infty$, we also have that $t^{\star}:=\max_{i\in S^{\star}}t_{i}<\infty.$ Our proof is complete after noting that for all $t\geq\max\{s^{\star},t^{\star}\}$, we have that $i_{t}\in S^{\star}$ and $\{z^{i_{t}}_{1},z^{i_{t}}_{2},\dots\}\setminus\{x_{1},\dots x_{t}\}\subseteq\operatorname{supp}(h).$
4 Discussion and Open Questions
In this paper, we introduced several notions of noisy generatability. We highlight some important directions of future work that we left open.
Characterizations of Noisy Non-uniform Generatability and Noisy Generatability in the Limit. An important direction of future work is to provide a complete and concise characterization of which classes are noisily non-uniformly generatable and which classes are noisily generatable in the limit.
Question 4.1.
What characterizes noisy non-uniform generatability and noisy generatability in the limit?
Note that Li et al. (2024) also pose the complete characterization of (noiseless) generatability in the limit as an open question, but do provide a complete characterization of (noiseless) non-uniform generatability. We conjecture that the correct characterization of noisy non-uniform generatability will need to go beyond our sufficiency and necessity conditions in Section 3.3.
Noisy Generatability in the Limit via Membership Oracles. In addition to showing that all countable classes are generatable in the limit, Kleinberg and Mullainathan (2024) provide a computable algorithm for doing so when given access to a membership oracle. In our notion, a membership oracle $\mathcal{O}_{\mathcal{H}}:\mathcal{H}\times\mathcal{X}\rightarrow\{0,1\}$, takes as input a hypothesis $h\in\mathcal{H}$ and an example $x\in\mathcal{X}$ and outputs $\mathbbm{1}\{h(x)=1\}$. Although we show that all countable classes are also noisily generatable in the limit, the focus of this paper was mainly information-theoretic in nature. This motivates our next open question.
Question 4.2.
Given access to a membership oracle, does there exists a computable algorithm that can noisily generate in the limit from any countable class?
Noiseless vs. Noisy Generatability In Section 3.2, we showed that noiseless uniform generatability can be strictly easier than noisy uniform generatability. Our third open question asks whether this is also true for non-uniform generatability.
Question 4.3.
Is there a class which is (noiseless) non-uniformly generatable but not noisily non-uniformly generatable?
Recall that in Section 3.3, we showed that all countable classes are noisily non-uniformly generatable. Thus, an affirmative resolution to Question 4.3 must involve an uncountably infinite class. Nevertheless, we conjecture that such a separation is true given the separation between noisy and noiseless uniform generation. One can also ask this question with regards to generatability in the limit.
Question 4.4.
Is there a class which is (noiseless) generatable in the limit but not noisily generatable in the limit?
Again, a positive affirmation to this questions must involve an uncountably infinite class.
Appendix A The Hardness of Non-uniform Noisy Generatability
As alluded to in Remark 2.3, there is a fifth setting of noisy generation, which we did not define in the main text. In this setting, the “sample complexity” of the generator can be non-uniform over the class $\mathcal{H}$ but must still be uniform over the noise-level and stream picked by the adversary. We define this formally below.
Definition A.1 (Non-uniform Noisy Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly noisily generatable, if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}\cap\operatorname{supp}(h)|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Unfortunately, non-uniform noisy generatability is still a restrictive property as not all finite classes are non-uniformly noisily generatable. This is in contrast to noisy uniform generatability as shown by Corollary 3.3.
Lemma A.1 (Not All Finite Classes are Non-uniformly Noisily Generatable).
Let $\mathcal{X}$ be countable. There exists a finite class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ which satisfies the UUS property that is not non-uniformly noisily generatable.
Proof.
Let $\mathcal{X}=\mathbbm{N}$ and consider the class =$\mathcal{H}=\{h_{e},h_{o}\}$ such that $h_{e}(x)=\mathbbm{1}\{x\text{ is even}\}$ and $h_{o}(x)=\mathbbm{1}\{x\text{ is odd}\}.$ We will show that $\mathcal{H}$ is not non-uniformly noisily generatable. We need to show that for every generator $\mathcal{G}$, there exists a hypothesis $h\in\mathcal{H}$ such that for every $d\in\mathbb{N}$, there exists a sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$
so that for every $t\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|=d$, there exists a $s\geq t$ where $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$. To that end, fix a generator $\mathcal{G}.$ Let $S=\{\mathcal{G}(1,2,\dots,i)\}_{i\in\mathbb{N}}.$ There are three cases to consider: (1) the number of even numbers in $S$ is finite (2) the number of odd numbers in $S$ is finite and (3) there are infinite number of both even and odd numbers in $S$. Cases (1) and (2) are symmetric. So, without loss of generality, it suffices only to consider Cases (1) and (3).
Starting with Case (1), let $p\in\mathbb{N}$ be the smallest time point such that for all $t\geq p$, we have that $\mathcal{G}(1,\dots,t)$ is odd. We will pick $h_{e}$ to use for the lower bound. Fix some $d\in\mathbb{N}$. Consider the sequence $x_{1},x_{2},\dots$ such that $x_{i}=i$ for all $i\leq\max\{p,2d\}$ and $x_{i}=2i$ for all $i>\max\{p,2d\}.$ Note that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h_{e})\}\leq\frac{\max\{p,2d\}}{2}<\infty.$
Suppose $2d<p.$ Then, observe that only on time point $t=2d$ do we have that |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h_{e})|=d$. However, by definition, on time point $s=p\geq 2d=t$, we have that $\mathcal{G}(x_{1},\dots,x_{s})$ is odd and therefore not in $\operatorname{supp}(h_{e}).$ Suppose that $2d\geq p.$ Then, like before, only on time point $t=2d$ do we have that |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h_{e})|=d$. However, by definition, also on time point $s=2d\geq p$, we have that $\mathcal{G}(x_{1},\dots,x_{s})$ is odd and therefore not in $\operatorname{supp}(h_{e}).$ Since $d\in\mathbb{N}$ is arbitrary, this is true for all $d\in\mathbb{N}$, completing the proof for Case (1). As mentioned before, the proof for Case (2) follows by symmetry.
We now consider Case (3). For this case, we will pick $h_{e}$ for the lower bound. Fix some $d\in\mathbb{N}$. Let $p\geq 2d$ be the smallest time point after and including time point $2d$ such that $\mathcal{G}(1,2,\dots,p)$ is odd. Such a $p\in\mathbb{N}$ must exist since $S$ contains an infinite number of odd numbers. Now, consider the sequence $x_{1},x_{2},\dots$ such that $x_{i}=i$ for $i\leq p$ and $x_{i}=2i$ for all $i>p.$ Note that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h_{e})\}\leq\frac{p}{2}<\infty.$
Observe that only at time point $t=2d$ do we have that |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h_{e})|=d.$ However, by definition, also on time point $s=p\geq 2d=t$, we have that $\mathcal{G}(x_{1},\dots,x_{p})$ is odd and therefore does not lie in $\operatorname{supp}(h_{e}).$ Since $d\in\mathbb{N}$ is arbitrary, this is true for all $d\in\mathbb{N}$, completing the proof for Case (3) and the overall proof. ∎